OpenStudy (anonymous):

Amita had to do a multiplication. Instead of taking 35 as one of the multipliers, she took 53. As a result the product went up by 540. What is the new product?

OpenStudy (anonymous):

actual product shpuld be (540/53)*35

OpenStudy (anonymous):

Let the product of other factors be n (n x 35) + 540 = n x 53 540 = 53n - 35n 540 = 18n 540/18 = n 30 = n New product = 30 x 53 = 1590

OpenStudy (anonymous):

prashant i m not satisfy your answer harkirat giving exact solution to this